# Self-Consistent Stochastic Finite-Temperature Modelling: Ultracold Bose Gases with Local (s-wave) and Long-Range (Dipolar) Interactions

### v1

@@ II.1 Basic Hamiltonian – Effective Action
A general non-local interaction contribution, $U_{nl}(\mathbf{r}-\mathbf{r}^{\prime})$, can arise in many distinct scenarios, including – for example – dipolar interactions, Rydberg interactions, or van der Waals interactions: details of the particular nature of the arising interaction in each case can be found in Ref. [162]. The theory presented in this work is directly applicable, provided we can formally write down a matrix inversion for $U_{nl}$.
Although the more general final form of equations for a general potential are given explicitly in Appendix A, in the remainder of this paper we have chosen to state our results explicitly in terms of a long-range interaction written in the form of the usual long-range dipolar gas interatomic potential form. This is in order to connect here directly to the community of dipolar atomic condensates, for which the form of the interatomic potential is very well known. With that in mind, we set $U_{nl}(\mathbf{r}-\mathbf{r}^{\prime})\rightarrow U_{dd}(\mathbf{r}-\mathbf{r}^{\prime})$, or
@@
the full non-local potential of Eq. (2) can be recast into the form
$U(\mathbf{r}-\mathbf{r}^{\prime})=\left(g-\frac{C_{dd}}{3}\right)\delta(\mathbf{r}-\mathbf{r}^{\prime})-C_{dd}\frac{1}{4\pi}(\mathbf{n}\cdot\nabla)^{2}\left(\frac{1}{r}\right)\;.$ (5)
@@
Following on from our detailed cosmological work in the context of dark matter [107], we consider the corresponding (non-relativistic) action for a bosonic particle $\psi(t,\mathbf{r})$ with mass $m$, defined as ($c=1$)
@@
Via a Hubbard-Stratonovich transformation we introduce an auxiliary field $V$ and recast the above Hamiltonian, mapping our problem to that of an equivalent effective action
@@
Apart from the time derivative term, the object in brackets is effectively equivalent to the hamiltonian in (6). To see this, it is enough to take the equation of motion for $V$ (obtained by varying the action (8) with respect to $V$) and plug it back into the action, resulting in the original hamiltonian. More details on this are given in Appendix A, where such steps are shown explicitly for a general operator.
Having recast the problem in terms of the above effective action (8), we can directly capitalize on our companion work deriving corresponding equations in the context of cosmological fuzzy dark matter where – in addition to local interactions – the particles are bound by the gravitational potential. To avoid duplication, and in the interest of clarity, we focus here directly on the implications of our model for ultracold atomic gases, without reproducing all the lengthy technical details of the derivation in this manuscript, referring instead the reader to such alternative work [106, 107]. Instead we only give here a brief overview of the key steps, highlighting the approximations made, which enable us to understand the limitations of our self-consistently obtained final set of equations.
II.2 A Summary of Key Steps:
@@
    with the first terms corresponding to the slow (or low-momentum) component, and the second part to the fast (or higher-momentum) part, noting that the location of such a split is arbitrary. Thus, the slow component $\Phi$ would typically contain elements with a higher order of coherence, but not (necessarily) fully condensed, relegating the incoherent part to $\varphi$. In the case of the auxiliary potential, we will work considering that $V_{d}$ and $V_{d}^{q}$ are slow quantities. Here we are ignoring possible fast contributions for the long-range potential, meaning that we are only taking in account (macroscopic) mean-field effects of this potential. Doing this, terms in the action with only one $\varphi$ term are discarded due to violation of energy-momentum conservation. This class of objects are: terms with one slow field $\Phi$ and one fast $\varphi$, terms with three slow $\Phi$ and one fast $\varphi$, and terms with one slow $\Phi$, one slow $V$ and one fast $\varphi$. Also, within this work, we will keep up to order two the components $\Phi^{q}$ and $V_{d}^{q}$, while simultaneously maintaining all orders of $\varphi$ (since we will integrate out this field).
@@
    We use the generating functional $Z=\int\mathcal{D}[\Phi V\varphi]e^{iS_{SK}}$, working in perturbation theory, expanding up to second order in the couplings related with the self interaction (i.e. $g-\frac{C_{dd}}{3}$) and the interaction with the auxiliary field $V$ (i.e. $C_{dd}$). Then, we integrate out the fast component $\varphi$ such that we are left with an effective action for the slow parts. In the process of integration we obtain lengthy expressions containing propagators and four-point functions of $\varphi$. In order to manage them, we make use of Wick’s Theorem and we discard ‘anomalous’ terms like $\langle\varphi\varphi\rangle$, $\langle\varphi^{\ast}\varphi^{\ast}\rangle$, $\langle\varphi^{q}\varphi^{q}\rangle$ and $\langle\varphi^{q\ast}\varphi^{q\ast}\rangle$ in what is often referred to as the ‘Popov’ approximation [56, 166, 167]. After integration, via a Hubbard-Stratonovich transformation we change the terms containing $(V_{d}^{q})^{2}$ and $(\Phi^{q})^{2}$ to objects linear in these fields by introducing auxiliary fields $\xi_{1}$ and $\xi_{2}$, which will be the noise terms appearing in our equations. Thus, we obtain an effective action for the slow fields, from which we compute the Euler-Lagrange equations after varying with respect to $\Phi^{q}$ and $V_{d}^{q}$, arriving at the equations of motion for the coherent part $\Phi_{0}$ and the field $V_{d}$ respectively.
@@
II.3 Relation to Companion Work (Ref. [107]):
@@
With this in mind, we can now directly write down the final equations relevant for the context of ultracold dipolar atomic gases. Although our emphasis here is on the experimentally-relevant cold atomic gases, we note that our mathematical approach is in fact more general, and can handle – at least in principle – any type of interaction potential, so our methodology might be useful to a broader class of systems. Details of such more general formulation, without resorting to any particular form of the interaction potential, are presented in Appendix A for any non-local interaction which can be cast in the form of an inverse operator $\cal{O}$(${}^{-1}({\bf r},\,{\bf r^{\prime}})$, as such formulation enables us to directly import our developed methodology to such a setting.
III Generalized Self-Consistent Stochastic Quantum Modelling for Dipolar Gases
Our final set of self-consistently derived equations is composed of
  • (i) A stochastic Gross-Pitaevskii equation for the classical, or ‘coherent’ field $\Phi_{0}$
@@ III Generalized Self-Consistent Stochastic Quantum Modelling for Dipolar Gases
    $\hskip 42.67912pt-2\left(g-\frac{C_{dd}}{3}\right)\int d^{4}x^{\prime}\Pi^{R}(x^{\prime},x)V_{nc}(x^{\prime})\Phi_{0}(x)+\left(g-\frac{C_{dd}}{3}\right)\Phi_{0}\,\xi_{2}$ (14)
    describing the low-lying highly-populated modes of the system, and featuring two collisional terms (identified by the ‘dissipative’ $R$ and ‘scattering’ $\Pi^{R}$) and their two corresponding complex additive ($\xi_{1}$) and real multiplicative ($\xi_{2}$) stochastic contributions [defined respectively in Eqs. (23), (24), (25), (26)].
  • (ii) A quantum Boltzmann equation for the high-lying thermal, or ‘incoherent’ modes
@@
  • (iii) A Poisson-like equation defining the effective dipolar potential $V_{d}$ in the form
@@
The 3 lines in the dynamical classical field equation (12) for $\Phi_{0}$ amount to a pure mean field dynamics term and two elastic collisional processes respectively, namely (i) an $O(g^{2})$ dissipative contribution facilitating particle transfer between the coherent (classical field) modes and the incoherent (high-lying) modes described by the Boltzmann equation, with its corresponding noise term $\xi_{1}$, and (ii) an $O(g^{2})$ scattering process between a coherent-band and an incoherent-band particle which however only leads to particle redistribution within each band and energy exchange between the bands, but not to direct particle transfer between those two bands, alongside its corresponding noise contribution. The meaning of these terms is further discussed within the context of the stochastic projected GPE (SPGPE) model of Sec. IV.3.2, a well-studied model where such terms first appeared, which emerges as a limiting case of our present theory.
@@
At this point it is pertinent to note that, as our derivation considers fluctuations of the fast $\varphi$-modes only, it does not self-consistently generate [^4] We note that while such a term can be routinely generated by adding a further term of the form $(2/5)\int d^{3}{\bf r}\gamma_{QF}|\Phi_{0}|^{5}$ to our energy functional [168], i.e. considering a more general effective action than that of the usual $\phi^{4}$ theory, we have not done this here as such a term is not present in the original, Hamiltonian and should arise as a correction to the classical, fluctuation-less dynamics when fluctuations of the relevant modes are integrated out. the correction term arising from the full inclusion of quantum fluctuations, known as the Lee-Huang-Yang (LHY) contribution [58, 169, 170, 171, 172, 173, 174]. Although ignoring such a term is reasonable in the rather dilute regime $na^{3}\ll 1$, long-range interactions can, under certain experimentally-accessible regimes, enhance the importance of this term. Following on from recent trends in the literature, and to connect with what is often being done, one could additionally include a term of the form $\gamma_{QF}|\Phi_{0}|^{3}\Phi_{0}$ to the RHS of Eq. (12), where $\gamma_{QF}$ is a function of the diluteness parameter $na^{3}$ [169, 170, 171, 172, 175, 176] which, more generally, is also temperature-dependent [177, 178, 168, 173]. Such a term is known to play a key role when mean-field interactions are suppressed such that quantum fluctuations become the dominant contribution, which has been well-studied in the context of stabilization of self-bound droplets [179, 180], with the inclusion of such a term into the usual GPE equation having become the norm in the field of dipolar condensates in recent years (A more extended list of references can be found in the recent review [162]): Although such a contribution to the equations does not emerge self-consistently within our treatment, instead requiring some additional quasiparticle handling, we will also comment on this further in Sec. V and Appendix C.
The above equations (12)-(14), (15) and (16) are general equations that can reduce to certain known limits and they can be connected with other formalisms in the literature. Here, we consider the well-studied case of typical ultracold quantum gases, interacting only via local effective interactions (s-wave scattering, Sec. IV), and the case of dipolar interactions (Sec. V), and discuss emerging common theories as limiting cases.
IV Reductions to Known Models for Locally-Interacting Atomic Gases
@@
IV.1 Mean Field Limit: The Gross-Pitaevskii equation
@@
Starting from Eq. (29), we can also include excitations of the condensate via a linearised treatment around $\Phi_{0}$, leading to the well-known Bogoliubov-de Gennes equations [8, 9, 181, 10, 19, 11, 12], containing quasiparticle physics. In the dilute regime $na^{3}\ll 1$, such contributions are generally small. However, as we shall see later, this point will become very relevant in the subsequent discussion of Sec. V in the context of dipolar gases.
IV.2 The Zaremba-Nikuni-Griffin (or ZNG) Model:
@@ IV.3 Stochastic models
IV.3.1 Stochastic Approach of Stoof

### v2

@@ III.1 Basic Hamiltonian – Effective Action
A general non-local interaction contribution, $U_{nl}(\mathbf{r}-\mathbf{r}^{\prime})$, can arise in many distinct scenarios, including – for example – dipolar interactions, or Rydberg interactions, with details of the particular nature of the arising interaction in each case found in Ref. [1]. Other systems where similar non-local interactions are relevant, include indirect exciton studies (see, e.g., [223]), or dipolar molecules [221, 224]. The theory presented in this work is directly applicable for any type of long-range interaction of the form of Eq. (2), provided we can formally write down a matrix inversion for $U_{nl}$ (see Appendix A).
Although the more general final form of equations for a general potential are given explicitly in Appendix A, in the remainder of this paper we have chosen to state our results explicitly in terms of a long-range interaction written in the form of the usual long-range dipolar gas interatomic potential form [225, 226, 227, 228, 229]. This is in order to connect here directly to the community of dipolar atomic condensates, for which the form of the interatomic potential is very well known. With that in mind, we set $U_{nl}(\mathbf{r}-\mathbf{r}^{\prime})\rightarrow U_{dd}(\mathbf{r}-\mathbf{r}^{\prime})$, or
@@
the full non-local potential of Eq. (2) can be recast into the form [230, 231, 232, 233]
$U(\mathbf{r}-\mathbf{r}^{\prime})=\left(g-\frac{C_{dd}}{3}\right)\delta(\mathbf{r}-\mathbf{r}^{\prime})-C_{dd}\frac{1}{4\pi}(\mathbf{n}\cdot\nabla)^{2}\left(\frac{1}{r}\right)\;,$ (5)
where one part of the dipolar interaction plays the role of renormalizing the strength of the local interaction.
@@
In order to proceed with our derivation, we now consider [^5] This procedure is motivated by our closely related cosmological work [165, 166], which derived a general set of self-consistent equations focussing on a gravitational (instead of dipolar) long-range interaction. the corresponding action for a bosonic particle $\psi(t,\mathbf{r})$ with mass $m$, defined as ($c=1$)
@@
Via a Hubbard-Stratonovich transformation we introduce an auxiliary field $V$ and recast the above Hamiltonian, mapping our problem to that of an equivalent effective action [^6] Note that we have chosen to apply the Hubbard-Stratonovich transformation to the so-called direct channel; the possible ramifications of a different choice [234], in conjunction with an appropriate fast-slow field split (discussed below) are left to future work. We thank Axel Pelster for raising the point with us.
@@
Apart from the time derivative term, the object in brackets is effectively equivalent to the hamiltonian in (6). To see this, it is enough to take the equation of motion for $V$ (obtained by varying the action (8) with respect to $V$) and plug it back into the action, resulting in the original hamiltonian. More details of such equivalence are given in Appendix A, cast (for generality) in the more general context of a general operator.
Having recast the problem in terms of the above effective action (8), we can now directly capitalize on our companion works [165, 166] deriving corresponding equations in the context of locally-interacting bosonic particles additionally bound by the gravitational potential, which plays the role of a long-range interaction. To avoid duplication, and in the interest of clarity, we focus here directly on the implications of our model for ultracold atomic gases, without reproducing all the lengthy technical details of the derivation in this manuscript, referring instead the reader to those works [165, 166]. Instead we only give here a brief overview of the key steps, highlighting the approximations made, which enable us to understand the limitations of our self-consistently obtained final set of equations.
III.2 A Summary of Key Steps:
@@
    with the first terms corresponding to the slow (or low-momentum) component, and the second part to the fast (or higher-momentum) part, noting that the location of such a split is arbitrary. Thus, the slow component $\Phi$ would typically contain elements with a higher order of coherence, but not (necessarily) fully condensed, relegating the incoherent part to $\varphi$. In the case of the auxiliary potential, we will work considering that $V_{d}$ and $V_{d}^{q}$ are slow quantities [^7] This appears to be a sensible prescription, although we note that it is not a priori guaranteed to correctly encompass all relevant diagrams, something which will be checked in future work.. Here we are ignoring possible fast contributions for the long-range potential, meaning that we are only taking in account (macroscopic) mean-field effects of this potential. Doing this, terms in the action with only one $\varphi$ term are discarded due to violation of energy-momentum conservation. This class of objects are: terms with one slow field $\Phi$ and one fast $\varphi$, terms with three slow $\Phi$ and one fast $\varphi$, and terms with one slow $\Phi$, one slow $V$ and one fast $\varphi$. Also, within this work, we will keep up to order two the components $\Phi^{q}$ and $V_{d}^{q}$, while simultaneously maintaining all orders of $\varphi$ (since we will integrate out this field).
@@
    We use the generating functional $Z=\int\mathcal{D}[\Phi V\varphi]e^{iS_{SK}}$, working in perturbation theory, expanding up to second order in the couplings related with the self interaction (i.e. $g-\frac{C_{dd}}{3}$) and the interaction with the auxiliary field $V$ (i.e. $C_{dd}$). Then, we integrate out the fast component $\varphi$ such that we are left with an effective action for the slow parts. In the process of integration we obtain lengthy expressions containing propagators and four-point functions of $\varphi$. In order to manage them, we make use of Wick’s Theorem and we discard [^8] Discarding such terms at this stage is equivalent to setting the coherence factors $u_{k}=1$ and $v_{k}=0$ in the Bogoliubov transformation, as discussed in [165]. See, however, Sec.VI for an alternative way of including such contributions and their potential relevance for dipolar gases. ‘anomalous’ terms like $\langle\varphi\varphi\rangle$, $\langle\varphi^{\ast}\varphi^{\ast}\rangle$, $\langle\varphi^{q}\varphi^{q}\rangle$ and $\langle\varphi^{q\ast}\varphi^{q\ast}\rangle$ in a manner closely related to the ‘Popov’ approximation [115, 235, 236]. After integration, via a Hubbard-Stratonovich transformation we change the terms containing $(V_{d}^{q})^{2}$ and $(\Phi^{q})^{2}$ to objects linear in these fields by introducing auxiliary fields $\xi_{1}$ and $\xi_{2}$, which will be the noise terms appearing in our equations. Thus, we obtain an effective action for the slow fields, from which we compute the Euler-Lagrange equations after varying with respect to $\Phi^{q}$ and $V_{d}^{q}$, arriving at the equations of motion for the coherent part $\Phi_{0}$ and the field $V_{d}$ respectively.
@@
  5. 5.
    Formulation in terms of dipolar interactions:
@@
    With this in mind, we can now directly write down the final equations relevant for the context of ultracold dipolar atomic gases.
    Although our emphasis here is on the experimentally-relevant cold atomic gases, we note that our mathematical approach is in fact more general, and can handle – at least in principle – any type of interaction potential, so our methodology might be useful to a broader class of systems. Details of such more general formulation, without resorting to any particular form of the interaction potential, are presented in Appendix A for any non-local interaction which can be cast in the form of an inverse operator $\cal{O}$(${}^{-1}({\bf r},\,{\bf r^{\prime}})$, as such formulation enables us to directly import our developed methodology to such a setting.
IV Generalized Self-Consistent Stochastic Quantum Modelling for Dipolar Gases
We can now directly write down the final set of equations arising through the above procedure. They include an equation for the low-lying modes of the system, $\Phi_{0}$, self-consistently coupled to a quantum Boltzmann equation for the high-lying thermal modes, and a Poisson-like equation for the (main) non-local part of the interaction, constituting (as we show below) a direct hybridization and extension of classical field and kinetic methods extensively used for modelling ultracold atomic gases interacting via contact interactions.
More specifically:
  • (i) The classical, or ‘coherent’, field $\Phi_{0}$ obeys a stochastic Gross-Pitaevskii equation in the form
@@ IV Generalized Self-Consistent Stochastic Quantum Modelling for Dipolar Gases
    $\hskip 42.67912pt-2\left(g-\frac{C_{dd}}{3}\right)\int d^{4}x^{\prime}\Pi^{R}(x^{\prime},x)V_{nc}(x^{\prime})\Phi_{0}(x)+\left(g-\frac{C_{dd}}{3}\right)\Phi_{0}\,\xi_{2}\,,$ (14)
    describing the low-lying highly-populated modes of the system. This equation features two collisional terms (identified by the ‘dissipative’ $R$ and ‘scattering’ $\Pi^{R}$) and their two corresponding complex additive ($\xi_{1}$) and real multiplicative ($\xi_{2}$) stochastic contributions [defined respectively in subsequent Eqs. (24), (25), (26), (27)]. Note that theoretical studies of dipolar condensates typically also include (particularly near the instability threshold) an additional term to the usual Gross-Pitaevskii equation [Eq. (12)] in the form
    $i\frac{\partial\Phi_{0}}{\partial t}=\cdots+\gamma_{QF}|\Phi_{0}|^{3}\Phi_{0}\;,$ (15)
    where $\gamma_{QF}$ [defined subsequently in Eq. (69)] is a function of the diluteness parameter $na^{3}$ [23, 24, 22, 25, 28, 27] which, more generally, is also temperature-dependent [42, 43, 44, 49]. Such a contribution, known as the Lee-Huang-Yang (LHY) contribution [117, 23, 24, 22, 25, 49, 237] provides a correction relevant for reproducing the ground state system energy. Although it is generally reasonable to ignore such a term for gases interacting via weak effective local (contact) interactions in the rather dilute regime $na^{3}\ll 1$, long-range interactions can, under certain experimentally-accessible regimes, enhance the importance of this term. Such a term is known to play a key role when mean-field interactions are suppressed such that quantum fluctuations become the dominant contribution, which has been well-studied in the context of stabilization of self-bound droplets [29, 8], with the inclusion of such a term into the usual GPE equation having become the norm in the field of dipolar condensates in recent years (A more extended list of references can be found in the recent review [1]). This contribution arises naturally upon linearizing the condensate wavefunction by means of the usual Bogoliubov-de Gennes analysis, but does not directly emerge from our above analysis due to our choice of setting $u_{k}=1$ and $v_{k}=0$ discussed in the previous section. Despite the significant success of such ‘quantum’ term in modelling droplets and supersolids in the $T=0$ limit [20], not all experiments necessarily require such extra contribution [14] (see also Appendix F in [59]), whose role in fact appears to decrease with increasing temperature [37, 59]. We defer all further discussion of this term, and its inclusion in phenomenological stochastic studies [57] to Sec. VI and Appendix B, with Eq. (77) giving a plausible criterion for its relative importance in a finite-temperature dipolar gas.
  • (ii) The high-lying thermal, or ‘incoherent’ modes obey a quantum Boltzmann equation in the form
@@
  • (iii) Non-local interactions are included through a Poisson-like equation defining the effective dipolar potential $V_{d}$ in the form
@@
The 3 lines in the dynamical classical field equation (12)-(14) for $\Phi_{0}$ amount respectively to a pure mean field dynamics term [(12)], and two elastic collisional processes respectively, namely (i) an $O(g^{2})$ dissipative contribution facilitating particle transfer between the coherent (classical field) modes and the incoherent (high-lying) modes described by the Boltzmann equation, with its corresponding noise term $\xi_{1}$ [(13)] , and (ii) an $O(g^{2})$ scattering process between a coherent-band and an incoherent-band particle which however only leads to particle redistribution within each band and energy exchange between the bands, but not to direct particle transfer between those two bands, alongside its corresponding noise contribution [(14)] . The meaning of these terms is further discussed within the context of the stochastic projected GPE (SPGPE) model of Sec. V.3.2, a well-studied model where such terms first appeared, which emerges as a limiting case of our present theory [^9] Note that the SPGPE model also makes explicit statements regarding the location of the cutoff relevant for its numerical implementations, with such aspects not directly addressed within our present work..
@@
The above equations (12)-(17) are general equations that can reduce to certain known limits and they can be connected with other formalisms in the literature. For clarity, we highlight once more that as our finite temperature derivation based on [166] considers fluctuations of the fast $\varphi$-modes only under the assumption of $(u_{k}=1\,,v_{k}=0)$, it does not self-consistently generate [^10] We note that while the LHY term can be routinely generated in a typical density-functional manner by adding a further term of the form $(2/5)\int d^{3}{\bf r}\gamma_{QF}|\Phi_{0}|^{5}$ to our energy functional [44], i.e. considering a more general effective action than that of the usual $\phi^{4}$ theory, we have not done this here as such a term is not explicitly additionally present in the original Hamiltonian, but its effect should emerge from the full Hamiltonian by appropriate handling (full Bogoliubov diagonalization). the LHY correction term [Eq. (15)].
In the remainder of this paper, we first consider the well-studied case of typical ultracold quantum gases, interacting only via local effective interactions (s-wave scattering, Sec. V), and then focus on the case of dipolar interactions (Sec. VI), discussing common theories emerging as limiting cases.
V Reductions to Known Models for Locally-Interacting Atomic Gases
@@
V.1 Mean Field Limit: The Gross-Pitaevskii equation
@@
Starting from Eq. (30), we can also include excitations of the condensate via a linearised treatment around $\Phi_{0}$, leading to the well-known Bogoliubov-de Gennes equations [67, 68, 238, 69, 79, 70, 71], containing quasiparticle physics. In the dilute regime $na^{3}\ll 1$, such contributions are generally small. However, as we shall see later, this point will become very relevant in the subsequent discussion of Sec. VI in the context of dipolar gases.
V.2 The Zaremba-Nikuni-Griffin (or ZNG) Model:
@@ V.3 Stochastic models
V.3.1 Stochastic Approach of Stoof